# Aligning Human Knowledge with Visual Concepts Towards Explainable Medical Image Classification

### v1

@@ Abstract
Although explainability is essential in the clinical diagnosis, most deep learning models still function as black boxes without elucidating their decision-making process. In this study, we investigate the explainable model development that can mimic the decision-making process of human experts by fusing the domain knowledge of explicit diagnostic criteria. We introduce a simple yet effective framework, Explicd, towards Explainable language-informed criteria-based diagnosis. Explicd initiates its process by querying domain knowledge from either large language models (LLMs) or human experts to establish diagnostic criteria across various concept axes (e.g., color, shape, texture, or specific patterns of diseases). By leveraging a pretrained vision-language model, Explicd injects these criteria into the embedding space as knowledge anchors, thereby facilitating the learning of corresponding visual concepts within medical images. The final diagnostic outcome is determined based on the similarity scores between the encoded visual concepts and the textual criteria embeddings. Through extensive evaluation of five medical image classification benchmarks, Explicd has demonstrated its inherent explainability and extends to improve classification performance compared to traditional black-box models.
@@ 1 Introduction
The advent of deep learning [8, 7] has profoundly transformed the field of medical image analysis [17, 5, 14] in lowering diagnostic costs [13] and improving diagnostic accuracy [2]. Despite these advancements, state-of-the-art deep neural networks often operate as black boxes. While they can achieve high performance in end-to-end image classification, they fail to provide the explainable rationale behind the decision-making process. This lack of transparency compromises the trust and validation by healthcare professionals, leading to potential errors and resistance of integrating AI-derived insights into clinical settings [6]. Unlike these black-box AI models, as illustrated in Fig. 1 (B), human experts make diagnoses by meticulously analyzing key image features from color, shape, size, or specific patterns regarding the disease symptom, staging, and outcome. In essence, human-expert decisions are grounded in a set of differential diagnostic criteria, enabling us to distinguish between various medical conditions with confidence and transparency. Developing explainable models that mirror the nuanced criteria defined by human knowledge is crucial to real-world AI clinical applications.
@@
Incorporating human knowledge into the medical image assessment faces a long-standing challenge in the cross-modal data integration. The surge of vision-language models (VLMs) [20, 9, 16] has focused on aligning image-text pairs in a unified representation space, opening up opportunities for the joint understanding of vision and language tasks. However, major general VLMs [20] are not trained on medical data. They often fall short when dealing with medical image-text pairs due to a significant shift in data distribution. Although efforts have been placed on VLMs [4, 26] pretrained on biomedical data, their performance still lags behind task-specific models in a broad range of medical image analysis tasks [24]. This is because they often struggle to align the associations between nuanced task-specific texts and the corresponding visual features. This inherent gap necessitates the development of better fine-grained alignment methods.
@@ 5 Discussion
To address the lack of transparency and interpretability in current deep learning models for medical image analysis, we proposed Explicd, a comprehensive framework that integrates diagnostic criteria queried from LLMs, aligning visual concepts towards explainable classification. Explicd offers a novel means to understanding diseases along human-understandable criteria axes. Our extensive experiments highlight Explicd’s superior performance over both traditional black-box approaches and existing explainable models, setting a new standard in both accuracy and interpretability. The clarity of Explicd’s decision-making process promises to bolster trust and facilitate the integration of AI in clinical diagnostics. Moving forward, we aim to expand the incorporation of broad human knowledge within our diagnostic criteria and to refine the hierarchical representations of visual concepts, allowing for a more nuanced exploration of disease diagnosis and management.

### v2

@@ Abstract
Although explainability is essential in the clinical diagnosis, most deep learning models still function as black boxes without elucidating their decision-making process. In this study, we investigate the explainable model development that can mimic the decision-making process of human experts by fusing the domain knowledge of explicit diagnostic criteria. We introduce a simple yet effective framework, Explicd, towards Explainable language-informed criteria-based diagnosis. Explicd initiates its process by querying domain knowledge from either large language models (LLMs) or human experts to establish diagnostic criteria across various concept axes (e.g., color, shape, texture, or specific patterns of diseases). By leveraging a pretrained vision-language model, Explicd injects these criteria into the embedding space as knowledge anchors, thereby facilitating the learning of corresponding visual concepts within medical images. The final diagnostic outcome is determined based on the similarity scores between the encoded visual concepts and the textual criteria embeddings. Through extensive evaluation of five medical image classification benchmarks, Explicd has demonstrated its inherent explainability and extends to improve classification performance compared to traditional black-box models. Code is available at https://github.com/yhygao/Explicd.
@@ 1 Introduction
The advent of deep learning [12, 7] has profoundly transformed the field of medical image analysis [21, 5, 18, 11] in lowering diagnostic costs [17, 9, 10] and improving diagnostic accuracy [26, 2]. Despite these advancements, state-of-the-art deep neural networks often operate as black boxes. While they can achieve high performance in end-to-end image classification, they fail to provide the explainable rationale behind the decision-making process. This lack of transparency compromises the trust and validation by healthcare professionals, leading to potential errors and resistance of integrating AI-derived insights into clinical settings [6]. Unlike these black-box AI models, as illustrated in Fig. 1 (B), human experts make diagnoses by meticulously analyzing key image features from color, shape, size, or specific patterns regarding the disease symptom, staging, and outcome. In essence, human-expert decisions are grounded in a set of differential diagnostic criteria, enabling us to distinguish between various medical conditions with confidence and transparency. Developing explainable models that mirror the nuanced criteria defined by human knowledge is crucial to real-world AI clinical applications.
@@
Incorporating human knowledge into the medical image assessment faces a long-standing challenge in the cross-modal data integration. The surge of vision-language models (VLMs) [24, 13, 20] has focused on aligning image-text pairs in a unified representation space, opening up opportunities for the joint understanding of vision and language tasks. However, major general VLMs [24] are not trained on medical data. They often fall short when dealing with medical image-text pairs due to a significant shift in data distribution [32]. Although efforts have been placed on VLMs [4, 31] pretrained on biomedical data, their performance still lags behind task-specific models in a broad range of medical image analysis tasks [29, 8]. This is because they often struggle to align the associations between nuanced task-specific texts and the corresponding visual features. This inherent gap necessitates the development of better fine-grained alignment methods.
@@ 5 Discussion
To address the lack of transparency and interpretability in current deep learning models for medical image analysis, we proposed Explicd, a comprehensive framework that integrates diagnostic criteria queried from LLMs, aligning visual concepts towards explainable classification. Explicd offers a novel means to understanding diseases along human-understandable criteria axes. Our extensive experiments highlight Explicd’s superior performance over both traditional black-box approaches and existing explainable models, setting a new standard in both accuracy and interpretability. The clarity of Explicd’s decision-making process promises to bolster trust and facilitate the integration of AI in clinical diagnostics. Moving forward, we aim to expand the incorporation of broad human knowledge within our diagnostic criteria and refine the hierarchical representations of visual concepts, allowing for a more nuanced exploration of disease diagnosis and management.
{credits}
5.0.1 Acknowledgements
This research has been partially funded by research grants to D. Metaxas through NSF: 2310966, 2235405, 2212301, 2003874, and FA9550-23-1-0417 and NIH 2R01HL127661.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.